OpenStudy (anonymous):

Finley Iggley's check register balance shows $1,959.75 at the time he receives his bank statement. His bank statement shows a balance of $2,827.00. Finley finds that check #525 for $600.00 and check #527 for $307.25 are still outstanding. He remembers that he took cash from an ATM, but never wrote down the amount. What is the most likely amount of Finley's cash withdrawal?

Directrix (directrix):

Bank statement shows a balance of $2,827.00. From that balance, subtract the sum of the two outstanding checks of ($600 + $307.25). Do that because those checks have already been written but have not yet been processed by the banks but will be because the money has been spent. $2,827.00 - ($600 + $307.25) = ? Post what you get, okay? After that, there is one additional step to get to the answer to the posted question. @Notgoodinmath

OpenStudy (anonymous):

40

Directrix (directrix):

Are you saying that $2,827.00 - ($600 + $307.25) = $40? If so, then that is not correct.

Directrix (directrix):

What is this: $2,827.00 - ($600 + $307.25) = ?

Directrix (directrix):

@Notgoodinmath